OpenStudy (anonymous):

Find the common difference of the arithmetic sequence shown. 1/4,,,3/8,,,1/2,,..... 1/16 1/8 1/4 , , , ...

OpenStudy (anonymous):

1/16?????

OpenStudy (dls):

\[\Huge \frac{3}{8}-\frac{1}{4}=\]

OpenStudy (anonymous):

I would say 5/8 but that doesn't seem to be a choice. Because if you make all of them with 8 on the bottom they seem to go up one 8th per step

OpenStudy (ybarrap):

We need to find d, the distance between each sequence: $$\frac{1}{4} + (n-1)d=a_n\text{ where, n=1,2,...}$$ So take $$a_n=\frac{3}{8}$$ $$\frac{1}{4}+(4-1)d=\frac{3}{8}\text{ then d = }\frac{1}{24}$$ We used n=4 because it's the 4th element in the sequence. So d=\(\dfrac{1}{24}\) is your answer.

OpenStudy (ybarrap):

You could have also used n=7 for \(a_7=\frac{1}{2}\) to find d, and would have found the same answer.

OpenStudy (ybarrap):

The formula for the arithmetic sequence I used above is documented here: http://en.wikipedia.org/wiki/Arithmetic_sequence